# Experimental Demonstration of a Decentralized Electromagnetic Formation Flying Control Using Alternating Magnetic Field Forces

### v1

@@ Abstract
Electromagnetic formation flying (EMFF) is challenging due to the complex coupling between the electromagnetic fields generated by each satellite in the formation. To address this challenge, this article uses alternating magnetic field forces (AMFF) to decouple the electromagnetic forces between each pair of satellites. Each satellite’s electromagnetic actuation system is driven by a sum of amplitude-modulated sinusoids, where amplitudes are controlled to achieve desired forces between each pair of satellites. The main contribution of this article is a 3-satellite experimental demonstration of decentralized closed-loop EMFF using AMFF. To our knowledge, this is the first demonstration of AMFF with at least 3 satellites in open or closed loop. This is noteworthy because the coupling challenges of EMFF are only present with more than 2 satellites, and thus, a formation of at least 3 is necessary to evaluate the effectiveness of AMFF. The experiments are conducted on a ground-based testbed consisting of 3 electromagnetically actuated satellites on linear air tracks. The closed-loop experimental results are compared with behavior from numerical simulations.
1 Introduction
Electromagnetic formation flying (EMFF) is accomplished using electromagnetic coils onboard satellites in a formation. Each satellite’s coils generate a magnetic field, which interacts with the magnetic fields of the other satellites to create magnetic field forces. These magnetic field forces can control the relative positions of satellites [10, 12, 14]. An advantage of EMFF is that the power source is renewable, while conventional actuation systems such as propellant thrusters eventually deplete.
EMFF for more than 2 satellites is challenging because the intersatellite forces are nonlinear functions of the magnetic moments generated by all satellites in the formation as well as the relative positions of all satellites. In other words, there is complex coupling between the electromagnetic fields generated by all satellites, and this coupling ultimately leads to intersatellite forces between every pair of satellites. EMFF control for more than 2 satellites is considered in [3, 15, 4]. However, these approaches require centralization of all measurement information and solving nonconvex constrained optimization problems that do not scale well to a large number of satellites.
A different approach to EMFF for more than 2 satellites is to use alternating magnetic field forces (AMFF) to address the challenge of coupling between the electromagnetic fields. The key idea of AMFF is that a pair of alternating (e.g., sinusoidal) magnetic moments results in a nonzero average interaction force between the pair of satellites if and only if those alternating magnetic moments have the same frequency [21, 13].
A decentralized EMFF method using AMFF is presented in [1]. In this method, each satellite has access to measurements of its position and velocity relative to only its local neighbor satellites. Then, each satellite uses a sum of frequency-multiplexed sinusoidal magnetic moments to achieve desired intersatellite forces between each pair of satellites. Other examples of EMFF using AMFF include [7, 22, 2, 18, 17, 19]. For example, [7, 22] uses AMFF for leader following. Relative position and attitude control with AMFF is addressed using a centralized algorithm in [19] and using a decentralized algorithm in [2]. In [18, 17], AMFF is combined with traditional reaction wheels in a centralized algorithm for relative position and attitude control.
Experimental evaluations of EMFF with AMFF are limited to [13, 2, 20]. Open-loop experiments demonstrating the AMFF approach are presented in [13]. Closed-loop single-degree-of-freedom experiments with a pair of satellites on linear air-tracks are presented in [2, 20]. Since the coupling challenges of EMFF are only present with more than 2 satellites, a formation of at least 3 satellites is necessary to evaluate the effectiveness of AMFF. However, to our knowledge, the existing literature does not include any experiments using AMFF with more than 2 satellites
The main contribution of this article is an experimental demonstration of closed-loop EMFF using AMFF with 3 satellites. To our knowledge, this is the first demonstration of AMFF with at least 3 satellites in open or closed loop. This work demonstrates EMFF with 3 satellites on a linear air track, where the coil of the center satellite is driven by a sum of 2 AC currents and each outer satellite is driven by one of the AC currents. As a result, intersatellite forces are generated between the center satellite and each outer satellite. This article adopts the piecewise-sinusoidal magnetic moment approach in [1] to address the challenges of coupling between the electromagnetic fields; this approach is reviewed in Section 4. Then, Section 5 presents the closed-loop decentralized EMFF algorithm, which is adopted from [1] but with a simplified approach for the control amplitude allocation. Section 6 describes the experimental testbed and modifies the EMFF control to account for hardware limitations such as sensor noise and amplitude saturation in the control current. Finally, experimental results are presented in Section 7. First, we present open- and closed-loop experiments with 2 satellites to demonstrate the feasibility of EMFF with AMFF. Finally, we present closed-loop experiments with 3 satellites, demonstrating frequency-multiplexed AMFF.
@@ 2 Notation
Let $\mathbb{N}$ denote the set of nonnegative integers. We let $\mathcal{I}\triangleq\{1,...,n\}$, where $n\in\mathbb{N}$ is the number of satellites in the formation, and $\mathcal{P}\triangleq\{(i,j)\in\mathcal{I}\times\mathcal{I}:i\neq j\}$, which is the set of ordered pairs. Unless otherwise stated, all statements in this paper that involve the subscript $i$, ijij, and $k$ are for all $i\in\mathcal{I}$, all $(i,j)\in\mathcal{P}$, and all $k\in\mathbb{N}$.
@@ 3 Dynamics
Each satellite has an electromagnetic actuation system (i.e., multiple electromagnetic coils) that creates a magnetic field. These magnetic fields interact to produce intersatellite forces which control the satellites’ relative positions. The coils are modeled as magnetic dipoles, and the resulting intersatellite force applied to satellite $i$ by satellite $j$ is
@@
where $\mu_{0}$ is the vacuum permeability constant, $\mathbf{u}_{i}$ is the magnetic moment of satellite $i$, and
@@
The magnetic moment $\mathbf{u}_{i}$, which is the control of satellite $i$, is a function of the current supplied to the electromagnetic coils. Since $\mathbf{r}_{ij}=-\mathbf{r}_{ji}$, it follows from (1) and (2) that $\mathbf{F}_{ij}=-\mathbf{F}_{ji}$, that is, the force applied to satellite $j$ by $i$ is equal and opposite the force applied to $i$ by $j$. See [3] and [15] for more information on the force model (1) and (2).
@@
where $c_{0}\triangleq 3\mu_{0}/(4\pi)$. Since $\mathbf{F}_{ij}=-\mathbf{F}_{ji}$, it follows from (3) that $\sum_{i\in\mathcal{I}}\dot{\mathbf{v}}_{i}(t)=0$, which implies that the linear momentum of the system of satellites is conserved. Thus, the intersatellite electromagnetic forces can be used to alter relative positions, but they have no effect on the overall mass center of the satellites.
Let the constant $\mathbf{d}_{ij}$ be the desired position of satellite $i$ relative to satellite $j$, where $\mathbf{d}_{ij}=-\mathbf{d}_{ji}$. The objective is to design controls that rely on feedback of neighboring satellites and yield $\lim_{t\to\infty}\mathbf{r}_{ij}(t)=\mathbf{d}_{ij}$.
4 Time-Averaged Dynamics with Piecewise-
Sinusoidal Controls
This section reviews the piecewise-sinusoidal magnetic-moment approach in [1], which is used to address the coupling that occurs between electromagnetic fields. In this approach, each satellite uses a piecewise-sinusoidal magnetic moment $\mathbf{u}_{i}$ that is the sum of n1n-1 sinusoids with n1n-1 unique frequencies. Each unique frequency is common to only one pair of satellites. Thus, there are a total of n(n1)/2n(n-1)/2 unique frequencies, and the amplitudes of each common-frequency pair of sinusoids are selected to prescribe the average intersatellite force between the associated satellite pair.
Let $\omega_{ij}\geq 0$ be the interaction frequency, where $\omega_{ij}=\omega_{ji}$ is unique. Next, let $T>0$ be a common multiple of $\{2\pi/\omega_{ij}:(i,j)\in{\mathcal{P}}\}$. For all $k\in\mathbb{N}$ and $t\in[kT,kT+T)$, we consider the piecewise-sinusoidal control
@@ 4 Time-Averaged Dynamics with Piecewise-
The following result characterizes the decoupling feature of the piecewise-sinusoidal control (4). This result is from [1, Proposition 1] and is a consequence of averaging the product of sinusoids over an integer number of periods.
@@
where the average intersatellite force is
@@
We define the approximate average intersatellite force
@@
which is an approximation of $\bar{\mathbf{F}}_{ij,k}$ obtained by replacing $\mathbf{r}_{ij}(t)$ with its sampling $\mathbf{r}_{ij,k}\triangleq\mathbf{r}_{ij}(kT)$. Propositon 1 implies that
@@
Note that $(\hat{\mathbf{F}}_{ij,k},\hat{\mathbf{F}}_{ji,k})$ depends on the amplitude control pair $(\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})$, which does not affect the approximate average intersatellite force between any other pair of satellites.

### v2

# arXiv:2601.05408v2[eess.SY] 02 Mar 2026

@@ Abstract
Electromagnetic formation flying (EMFF) is challenging due to the complex coupling between the electromagnetic fields generated by each satellite in the formation. To address this challenge, this article uses alternating magnetic field forces (AMFF) to decouple the electromagnetic forces between each pair of satellites. The key idea of AMFF is that a pair of alternating (e.g., sinusoidal) magnetic moments results in a nonzero time-averaged interaction force if and only if those alternating magnetic moments have the same frequency. Hence, the approach in this article is to drive each satellite’s electromagnetic actuation system with a sum of sinusoids, where each frequency is common to only a pair of satellites. Then, the amplitudes of each sinusoid are modulated (i.e., controlled) to achieve the desired forces between each pair of satellites. The main contribution of this article is a 3-satellite experimental demonstration of decentralized closed-loop EMFF using AMFF. To the authors’ knowledge, this is the first demonstration of AMFF with at least 3 satellites in open or closed loop. This is noteworthy because the coupling challenges of EMFF are only present with more than 2 satellites, and thus, a formation of at least 3 is necessary to evaluate the effectiveness of AMFF. The experiments are conducted on a ground-based testbed consisting of 3 electromagnetically actuated satellites on linear air tracks. The closed-loop experiments demonstrate decentralized EMFF with AMFF where the mean steady-state formation error is less than $0.005$ m, the maximum steady-state formation error is less than $\pm 0.01$ m, and the settling time is less than $30$ s. The closed-loop experimental results are compared with behavior from numerical simulations.
@@ 1 Notation
Let $\mathbb{N}$ denote the set of nonnegative integers. Let $\mathcal{I}\triangleq\{1,...,n\}$, where $n\in\mathbb{N}$ is the number of satellites in the formation, and $\mathcal{P}\triangleq\{(i,j)\in\mathcal{I}\times\mathcal{I}:i\neq j\}$, which is the set of ordered pairs. Unless otherwise stated, all statements in this paper that involve the subscript $i$, ijij, and $k$ are for all $i\in\mathcal{I}$, all $(i,j)\in\mathcal{P}$, and all $k\in\mathbb{N}$.
2 Nomenclature
$\mathcal{F}=\begin{bmatrix}\mathbf{i}&\mathbf{j}&\mathbf{j}\end{bmatrix}$ Inertial frame
$\mathbf{r}_{i}$ Position of satellite $i$ (m)
$\mathbf{v}_{i}$ Velocity of satellite $i$ (m/s)
$\mathbf{r}_{ij}\triangleq\mathbf{r}_{i}-\mathbf{r}_{j}$ Relative position (m)
$\mathbf{F}_{ij}$ Force applied to satellite $i$ by $j$ (N)
$\mathbf{f}$ Intersatellite force function (A2-m4)
$\mu_{0}$ Vacuum permeability constant (H/m)
$m$ Mass of each satellite (kg)
$\mathbf{u}_{i}$ Magnetic moment of satellite $i$ (A-m2)
$\mathbf{d}_{ij}$ Desired relative position (m)
$\omega_{ij}$ Interaction frequency (rad/s)
$\mathbf{p}_{ij,k}$ Amplitude control (A-m2)
$T$ Update period for control (s)
$\bar{\mathbf{F}}_{ij}$ Time-averaged intersatellite force (N)
$\hat{\mathbf{F}}_{ij}$ Approximate time-averaged
intersatellite force (N)
$\mathbf{r}_{ij,k}$ Sampled relative position
$\mathbf{v}_{ij,k}$ Sampled relative velocity (m/s)
$\mathbf{f}^{*}_{ij,k}$ Desired value of $\mathbf{f}(\mathbf{r}_{ij,k},\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})$ (A2-m4)
$\alpha_{ij}$, $\rho_{ij}$, $\beta$ Control gains
$r_{ij,k}$ Relative position in $\mathbf{i}$ direction (m)
$v_{ij,k}$ Relative velocity in $\mathbf{i}$ direction (m/s)
$d_{ij}$ Desired relative position in $\mathbf{i}$ direction (m)
$f^{*}_{ij,k}$ $\mathbf{f}^{*}_{ij,k}$ in $\mathbf{i}$ direction (A2-m4)
$p_{ij,k}$ Amplitude control in $\mathbf{i}$ direction (A-m2)
$I_{i}$ Current in the coil of satellite $i$ (A)
$I_{ij,k}$ Current amplitude (A)
$f$ $\mathbf{f}$ in $\mathbf{i}$ direction (A2-m4)
$\xi_{ij,k}$ Integrator state (m)
$\bar{I}$ Maximum allowable current (A)
$\hat{F}(r_{ij},p_{ij},p_{ji})$ $\hat{\mathbf{F}}_{ij}$ in $\mathbf{i}$ direction (N)
$q_{ij,k}$ Measurement of $r_{ij}$ (m)
$\hat{r}_{ij,k}$ Estimate of $r_{ij}$ (m)
$\hat{v}_{ij,k}$ Estimate of $v_{ij}$ (m/s)
$V$ Variance of noise associate with $q_{ij,k}$ (m2)
$L$ Kalman filter gain
$P$ Solution to algebraic Riccati equation
$w$ Variance of disturbance force (m2/s4)
$\hat{r}_{\mathrm{os}}$ Overshoot (m)
$T_{{\rm s}}$ Settling time (s)
$P_{\hat{F}}$ Root mean square of $\hat{F}$ (N)
3 Introduction
Electromagnetic formation flying (EMFF) is accomplished using electromagnetic coils that are onboard satellites in a formation. Each satellite’s coils generate a magnetic field, which interacts with the magnetic fields of the other satellites to create magnetic field forces. These magnetic field forces can control the relative positions of satellites [11, 13, 15]. An advantage of EMFF is that the power source is renewable, while conventional actuation systems such as propellant thrusters eventually deplete.
EMFF for more than 2 satellites is challenging because the intersatellite forces are nonlinear functions of the magnetic moments generated by all satellites as well as the relative positions of all satellites. In other words, there is complex coupling between the electromagnetic fields generated by all satellites, and this coupling ultimately leads to intersatellite forces between every pair of satellites. EMFF control for more than 2 satellites is considered in [4, 17, 5]. However, these approaches require centralization of all measurement information and solving nonconvex constrained optimization problems that do not scale well to a large number of satellites. An alternative approach in [16] relies on scheduling the actuation of each EMFF satellite over time. In this case, only a subset of the satellites are actuated for each time interval, which limits computational complexity because control computations are computed for only a subset of satellites. However, this approach requires centralization of the schedule and the control is computed through a nonconvex optimization. In sum, the coupling between electromagnetic fields causes scalability and decentralization challenges for EMFF.
A different approach to EMFF is to use alternating magnetic field forces (AMFF) to address the challenge of coupling between the electromagnetic fields. The key idea of AMFF is that a pair of alternating (e.g., sinusoidal) magnetic moments results in a nonzero average interaction force between a pair of satellites if and only if those alternating magnetic moments have the same frequency [23, 14].
A decentralized EMFF method using AMFF is presented in [1]. In this method, each satellite computes its desired intersatellite forces using measurements of its position and velocity relative to only its local neighbor satellites, avoiding the centralization and scalability limitations discussed above. Then, each satellite uses a sum of frequency-multiplexed sinusoidal magnetic moments to achieve desired intersatellite forces between each pair of satellites. For example, consider 3 satellites, where satellite 1 generates a magnetic moment that is the sum of sinusoids at 10 Hz and 20 Hz, satellite 2 generates a magnetic moment at 10 Hz, and satellite 3 generates a magnetic moment at 20 Hz. In this scenario, the average force between satellites 1 and 2 is a function of the amplitudes of the sinusoids at the common frequency 10 Hz. Notably, the average force between satellites 1 and 2 is independent of the sinusoidal amplitudes at 20 Hz. Similarly, the average force between satellites 1 and 3 is a function of the amplitudes at the common frequency 20 Hz. Finally, since the magnetic moments of satellites 2 and 3 do not have a common frequency, the average force between these satellites is negligible.
In addition to [1], EMFF using AMFF has been explored in [8, 24, 2, 20, 19, 21]. For example, [8, 24] presents a leader-following method with AMFF using sliding-mode control. Relative position and attitude control with AMFF is addressed with a centralized algorithm in [21] and using a decentralized algorithm in [2]. In [20, 19], AMFF is combined with traditional reaction wheels in a centralized algorithm for relative position and attitude control in the presence of disturbance torques caused by interaction with Earth’s magnetic field.
Experimental evaluations of EMFF with AMFF are in [14, 3, 22]. Open-loop experiments demonstrating the AMFF approach are presented in [14]. Closed-loop single-degree-of-freedom experiments with a pair of satellites on linear air-tracks are presented in [3, 22]. Note that since AMFF uses sinusoidal currents, it can be implemented with standard coils (as demonstrated in [14, 3, 22]) as opposed to superconducting coils, which have been used in DC-actuation experiments. Since the coupling challenges of EMFF are only present with more than 2 satellites, a formation of at least 3 satellites is necessary to evaluate the effectiveness of AMFF. However, to the authors’ knowledge, the existing literature does not include any experiments using AMFF with more than 2 satellites
The main contribution of this article is an experimental demonstration of closed-loop EMFF using AMFF with 3 satellites. This is the first demonstration of AMFF with at least 3 satellites in open or closed loop. This work demonstrates EMFF with 3 satellites on a linear air track, where the coil of the center satellite is driven by a sum of 2 AC currents and each outer satellite is driven by one of the AC currents. As a result, intersatellite forces are generated between the center satellite and each outer satellite. This article adopts the piecewise-sinusoidal magnetic moment approach in [1] to address the challenges of coupling between the electromagnetic fields; this approach is reviewed in Section 6. Then, Section 7 presents the closed-loop decentralized EMFF algorithm, which is adopted from [1] but with a simplified approach for the control amplitude allocation. Section 8 describes the experimental testbed and modifies the EMFF control to account for hardware limitations such as sensor noise and amplitude saturation in the control current. Experimental results are presented in Section 9. First, this article presents open- and closed-loop experiments with 2 satellites to demonstrate the feasibility of EMFF with AMFF. Next, this article presents closed-loop experiments with 3 satellites, where the satellites achieve formation in less than 30 s with mean steady-state error less than $0.005$ m and maximum steady-state formation error is less than $\pm 0.01$ m. This experiment demonstrates the decoupling principle of frequency-multiplexed AMFF.
4 Method Overview
The core concept of AMFF is that a pair of sinusoidal magnetic moments results in a nonzero time-averaged interaction force between a pair of satellites if and only if those alternating magnetic moments have the same frequency. The method in this paper is decentralized, which means that each satellite has access to measurements of its position and velocity relative to only its local neighbor satellites. First, this article presents a time-averaged model for the dynamics of a satellite formation, where the controllable magnetic moments are a sum of amplitude-modulated sinusoids. Next, the article shows how the sinusoidal amplitudes are selected to achieve desired intersatellite forces, which are determined decentrally by each satellite using feedback of its neighbors.
The main contribution of this article is the first experimental demonstration of EMFF using AMFF with more than 2 satellites. First, open- and closed-loop experiments are conducted with 2 satellites to demonstrate feasibility of AMFF. Next, closed loop experiments with 3 satellites are performed to demonstrate the core concept of AMFF.
@@ 5 Satellite Dynamics
Each satellite has an electromagnetic actuation system (i.e., multiple electromagnetic coils) that creates a magnetic field. These magnetic fields interact to produce intersatellite forces, which control the satellites’ relative positions. The coils are modeled as magnetic dipoles, and the resulting force applied to satellite $i$ by satellite $j$ is
@@
where $\mu_{0}$ is the vacuum permeability constant, $\mathbf{u}_{i}$ is the magnetic moment of satellite $i$, and the intersatellite force function is
@@
The magnetic moment $\mathbf{u}_{i}$, which is the control of satellite $i$, is a function of the current supplied to the electromagnetic coils. Since $\mathbf{r}_{ij}=-\mathbf{r}_{ji}$, it follows from 1 and 2 that $\mathbf{F}_{ij}=-\mathbf{F}_{ji}$, that is, the force applied to satellite $j$ by $i$ is equal and opposite the force applied to $i$ by $j$. See [4] and [17] for more information on 1 and 2 .
@@
where $c_{0}\triangleq 3\mu_{0}/(4\pi)$. Since $\mathbf{F}_{ij}=-\mathbf{F}_{ji}$, it follows from (3) that $\sum_{i\in\mathcal{I}}\dot{\mathbf{v}}_{i}(t)=0$, which implies that the linear momentum of the system of satellites is conserved. Thus, the intersatellite electromagnetic forces can be used to alter relative positions, but they have no effect on the overall mass center of the satellites. Therefore, traditional actuation systems may be needed in combination with EMFF to control the formation’s mass center. For example, propellant thrusters may be needed for orbit maintenance in the presence of disturbances (e.g., atmospheric drag in low Earth orbit) or to allow for orbit transfers. Similarly, a traditional attitude control system, such as reaction wheels, could be required for attitude control and addressing disturbances such as torques introduced by interactions with the Earth’s magnetic field.
Let the constant $\mathbf{d}_{ij}$ be the desired position of satellite $i$ relative to satellite $j$, where $\mathbf{d}_{ij}=-\mathbf{d}_{ji}$. The objective is to design and demonstrate a decentralized control method that relies on feedback of neighboring satellites and yields $\lim_{t\to\infty}\mathbf{r}_{ij}(t)=\mathbf{d}_{ij}$.
6 Time-Averaged Formation Dynamics with Sinusoidal Controls
This section reviews the piecewise-sinusoidal magnetic-moment approach in [1], which is used to address the coupling that occurs between electromagnetic fields. This approach enables decentralized formation control, where each satellite uses relative position and velocity measurements of only its neighbors. In this approach, each satellite uses a piecewise-sinusoidal magnetic moment $\mathbf{u}_{i}$ that is the sum of up to n1n-1 sinusoids with unique frequencies. Each unique frequency is common to only one pair of satellites. Thus, there are a total of up to n(n1)/2n(n-1)/2 unique frequencies, and the amplitudes of each common-frequency pair of sinusoids are selected to prescribe the time-averaged intersatellite force between the associated satellite pair.
Let $\omega_{ij}\geq 0$ be the interaction frequency between satellite $i$ and $j$, where $\omega_{ij}=\omega_{ji}$ is unique if it is nonzero. If an interaction frequency is zero, then the time-averaged intersatellite force between the associated satellite pair is zero. Next, let $T>0$ be a common multiple of $\{2\pi/\omega_{ij}:(i,j)\in{\mathcal{P}}\mbox{ and }\omega_{ij}\neq 0\}$. For all $k\in\mathbb{N}$ and $t\in[kT,kT+T)$, consider the piecewise-sinusoidal control
@@ 6 Time-Averaged Formation Dynamics with Sinusoidal Controls
Figure 2: AMFF with 3 satellites and unique frequencies $\omega_{12}=\omega_{21}$, $\omega_{13}=\omega_{31}$, and $\omega_{23}=\omega_{32}$. For each pair of satellites, the intersatellite forces ${\mathbf{F}}_{ij}=-{\mathbf{F}}_{ji}$ are determined on average from the amplitudes $(\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})$ at frequency $\omega_{ij}$.
The key idea of AMFF is that the amplitude pair $(\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})$ at the common frequency $\omega_{ij}$ determines the time-averaged forces between satellites $i$ and $j$. Figure 2 illustrates this concept for 3 satellites, where the components of $\mathbf{u}_{i}$ and $\mathbf{u}_{j}$ at $\omega_{ij}$ cause the time-averaged forces between satellites $i$ and $j$. The following result characterizes this decoupling property, which allows us to formalize the core concept of AMFF. The result is from [1, Proposition 1], and it is a consequence of averaging the product of sinusoids over an integer number of periods.
@@
where the time-averaged intersatellite force is
@@
We define the approximate time-averaged intersatellite force
@@
which is an approximation of $\bar{\mathbf{F}}_{ij,k}$ obtained by replacing $\mathbf{r}_{ij}(t)$ with its sampling $\mathbf{r}_{ij,k}\triangleq\mathbf{r}_{ij}(kT)$. Proposition 1 implies that
@@
Note that $(\hat{\mathbf{F}}_{ij,k},\hat{\mathbf{F}}_{ji,k})$ depends on the amplitude control pair $(\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})$, which does not affect the approximate time-averaged intersatellite force between any other pair of satellites. Thus, if $\mathbf{r}_{ij}$ does not change significantly over $[kT,kT+T)$, then the time-averaged formation dynamics can be approximated by (5) with $\bar{\mathbf{F}}_{ij}(k)$ replaced by $\hat{\mathbf{F}}_{ij,k}$.